OpenStudy (anonymous):

Given (triangle) JKL, with side lengths JK=22 and KL=30, what are the possible lengths of JL? Explain your reasoning

OpenStudy (amistre64):

the location of all points of a third vertex would form a circle around a stick: 

OpenStudy (amistre64):

these are the 2 legs, one rotating about the connecting point

OpenStudy (amistre64):

the shortest third leg possible is therefore the difference between the legs, and the longest possible is the sum of the legs. all other lengths range between these.

OpenStudy (anonymous):

So?

OpenStudy (anonymous):

its 13?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what is the difference in the legs? what is the sum of the legs?

OpenStudy (amistre64):

30-22=? 30+22=?

OpenStudy (anonymous):

13 and 52?

OpenStudy (anonymous):

oh okay 52 - 13?

OpenStudy (anonymous):

39

Directrix (directrix):

By the Triangle Inequality Theorem, one side of a triangle is less than the sum of the other two but greater than the positive difference. So (30 - 22) < JL < (22 + 30)

Directrix (directrix):

What is 30 - 22 =

OpenStudy (anonymous):

8

Directrix (directrix):

And, what is 22 + 30=

OpenStudy (anonymous):

52

Directrix (directrix):

8 < JL < 52

Directrix (directrix):

JL could have lengths between but not including 8 and 52

OpenStudy (anonymous):

Oh okay so lengths 9-51?

Directrix (directrix):

No

OpenStudy (anonymous):

?

Directrix (directrix):

The lengths were not specified to be integers. For all we know, JL could be 51.99

Directrix (directrix):

You are thinking in integers.

Directrix (directrix):

51 is not the greatest number less than 52. There is no such number.

Directrix (directrix):

JL could have lengths between but not including 8 and 52 which symbolically is this: 8 < JL < 52

Directrix (directrix):

Put this >>> okay so lengths 9-51? and you'll get it wrong.